# Identifying Functionally Important Features with End-to-End Sparse Dictionary Learning

### v1

@@ 8 Impact statement
This article proposes an improvement to methods used in mechanistic interpretability. Mechanistic interpretability, and interpretability broadly, promises to let us understand the inner workings of neural networks. This may be useful for debugging and improving issues with neural networks. For instance, it may be enable the evaluation of a model’s fairness or bias. Interpretability may relatedly be useful for improving the trust-worthiness of AI systems, potentially enabling AI’s use in certain high stakes settings, such as healthcare, finance, and justice. However, increasing the trust-worthiness of AI systems may be dual use in that may also enable its use in settings such as military applications.
@@ A.2 Pareto curves for TinyStories-1M
Note that most of our Tinystories-1M runs were for $\text{SAE}_{\text{local}}$ and $\text{SAE}_{\text{e2e}}$, and we did not perform several of the analyses that we performed for GPT2-small elsewhere in this report. But the clear improvement in $L_{0}$ and alive_dict_elements vs CE loss increase was apparent for $\text{SAE}_{\text{e2e}}$ vs $\text{SAE}_{\text{local}}$. More results can be found at https://api.wandb.ai/links/sparsify/yk5etolk.
@@ A.7 Auto-interpretability
Figure 9: Comparison of auto-interpretability scores between $\text{SAE}_{\text{e2e+ds}}$ and $\text{SAE}_{\text{local}}$. We choose two pairs at every layer, one with similar $L_{0}$ (see Table 3) and the other with similar CE loss increase (see Table 2). Error bars are a bootstraped 95% confidence interval for the true mean auto-interpretability scores. Measured on approximately $200(\pm 2)$ randomly selected features per dictionary.
Table 4: Estimates of the difference between the mean auto-interpretability scores for $\text{SAE}_{\text{e2e+ds}}$ and $\text{SAE}_{\text{local}}$ (Figure 9). A positive difference indicates $\text{SAE}_{\text{e2e+ds}}$ is more interpretable. For each comparison we use bootstrapping to compute a 95% confidence interval and a two-tailed p-value that the means are equal.
@@
In section 3.4 we claim that when comparing auto-interpretability scores we find no difference between pairs of similar $L_{0}$, but do find $\text{SAE}_{\text{e2e+ds}}$ is more interpretable than $\text{SAE}_{\text{local}}$ in layers 2 and 6. These results are presented in more detail in Figure 9 and Table 4.
@@ Appendix D Experimental details and hyperparameters
We stream the pre-tokenized dataset https://huggingface.co/datasets/apollo-research/Skylion007-openwebtext-tokenizer-gpt2 (tokenized as in GPT2 training, with a ‘BOS’ token between documents).
We evaluate our models on $500$ samples of Open Web Text (a different seed to that used for training). We consider a dictionary element alive if it activates at all on $500$k training tokens.
@@ Appendix H Training time
Table 6: Training times for different layers and SAE training methods using a single NVIDIA A100100100100 GPU on the residual stream of GPT2-small. All SAEs are trained on $400$k samples of context length $1024$, with a dictionary size of $60$x the residual stream size of $768$.

### v2

@@ 8 Impact statement
This article proposes an improvement to methods used in mechanistic interpretability. Mechanistic interpretability, and interpretability broadly, promises to let us understand the inner workings of neural networks. This may be useful for debugging and improving issues with neural networks. For instance, it may enable the evaluation of a model’s fairness or bias. Interpretability may relatedly be useful for improving the trust-worthiness of AI systems, potentially enabling AI’s use in certain high stakes settings, such as healthcare, finance, and justice. However, increasing the trust-worthiness of AI systems may be dual use in that may also enable its use in settings such as military applications.
@@ A.2 Pareto curves for TinyStories-1M
Note that most of our Tinystories-1M runs were for $\text{SAE}_{\text{local}}$ and $\text{SAE}_{\text{e2e}}$, and we did not perform several of the analyses that we performed for GPT2-small elsewhere in this report. But the clear improvement in $L_{0}$ and alive_dict_elements vs CE loss increase was apparent for $\text{SAE}_{\text{e2e}}$ vs $\text{SAE}_{\text{local}}$. More results can be found at https://api.wandb.ai/links/sparsify/yk5etolk. Future work would test that these results hold on more models of different sizes and architectures, as well as on SAEs trained not just on the residual stream.
@@ A.7 Automated interpretability
Figure 9: Comparison of automated interpretability scores between $\text{SAE}_{\text{e2e+ds}}$ and $\text{SAE}_{\text{local}}$. We choose two pairs at every layer, one with similar $L_{0}$ (see Table 3) and the other with similar CE loss increase (see Table 2). Error bars are a bootstraped 95% confidence interval for the true mean automated interpretability scores. Measured on approximately $200(\pm 2)$ randomly selected features per dictionary.
Table 4: Estimates of the difference between the mean automated interpretability scores for $\text{SAE}_{\text{e2e+ds}}$ and $\text{SAE}_{\text{local}}$ (Figure 9). A positive difference indicates $\text{SAE}_{\text{e2e+ds}}$ is more interpretable. For each comparison we use bootstrapping to compute a 95% confidence interval and a two-tailed p-value that the means are equal.
@@
In section 3.4 we claim that when comparing automated interpretability scores we find no difference between pairs of similar $L_{0}$, but do find $\text{SAE}_{\text{e2e+ds}}$ is more interpretable than $\text{SAE}_{\text{local}}$ in layers 2 and 6. These results are presented in more detail in Figure 9 and Table 4.
@@ Appendix D Experimental details and hyperparameters
For GPT2-small we stream the dataset https://huggingface.co/datasets/apollo-research/Skylion007-openwebtext-tokenizer-gpt2 which is a tokenized version of OpenWebText ([Gokaslan and Cohen, 2019]) (released under the license CC0-1.0). The tokenization process is the same as was used in GPT2 training, with a ‘BOS’ token between documents.
We evaluate our models on $500$ samples of the Open Web Text dataset (a different seed to that used for training). We consider a dictionary element alive if it activates at all on $500$k training tokens.
@@
We used NVIDIA A100100100100 GPUs with $80$GB VRAM (although the GPU was saturated when using smaller batch sizes that used $40$GB VRAM or less).
Our library imports from the TransformerLens library ([Nanda and Bloom, 2022], released under the MIT License) which is used to download models (among other things) via HuggingFace’s Transformers library ([Wolf et al., 2020], released under the Apache License 2.0). GPT2-small is released under the MIT license. The Tinystories-1M model is released under the Apache License 2.0 and it’s accompanying dataset is released under CDLA-Sharing-1.0.
@@ Appendix H Training time
Table 6: Training times for different layers and SAE training methods using a single NVIDIA A100100100100 GPU on the residual stream of GPT2-small at layer $6$. All SAEs are trained on $400$k samples of context length $1024$, with a dictionary size of $60$x the residual stream size of $768$.